A sample of helium has a volume of 34.8 L at a temperature of 25.0˚C. What would the volume be if the temperature was increased to 60.0˚C at a constant pressure?


1
Expert's answer
2021-04-07T11:30:07-0400

"\frac{V_1}{T_1}=\frac{V_2}{T_2}"


New volume is "V_2"


"V_2= \frac{V_1×T_2}{T_1}"


"V_2=\frac{34.8×[273+60]}{[273+25]}=38.89L"